OpenStudy (anonymous):

find the value of x (x+10)=(4x-35)

OpenStudy (rmrjr22):

x-4x=-35-10

OpenStudy (mathstudent55):

The first step is to drop the parentheses bec they are not needed in this equation.

OpenStudy (mathstudent55):

Then subtract 4x from both sides. Then subtract 10 from both sides.

OpenStudy (luigi0210):

Or do the opposite: 10-35=4x-x

OpenStudy (anonymous):

so im trying to get x by itself?

OpenStudy (luigi0210):

Yup!

OpenStudy (anonymous):

k but if i subtract ten from both sides it stil wont make x a loner... it wil make it -9.6

OpenStudy (lopezking1):

x=15

OpenStudy (anonymous):

x + 10 = 4x - 35 (subtract 4x from each side) x + 10 - 4x = 4x - 4x - 35 (combine like terms) -3x + 10 = - 35 do you understand so far ?

OpenStudy (luigi0210):